Let's assume that a firm's total weekly costs are as follows: 1. Salaries = $5000. 2. Supplies = $1000. 3. Rent = $600. In addition, the owners have invested $30,000 of their own money into the business. This could have earned them interest of $100 per week if they had chosen to put it into a bank instead of investing it into their business. If the firm has weekly revenue of $20,000, the firm's accounting profit is ____________, and the economic profit is _____________.

  

Loss of $18,987; profit of $5,690

   

Loss of $16,600; loss of $16,700.

   

Loss of $16,700; loss of $16,600.

   

$13,300; $13,400.

   

$13,400; $13,300.
Answer #1

Accounting Profit = Total Revenue - Total cost i.e. Explicit cost

= 20,000 - (Salaries + Rent + Supplies) = 20000 - (5000 + 600 + 1000) = 20000 - 6600 = 13400

Economic Profit = Total revenue - Total cost i.e. Implicit and Explicit cost

= 20,000 - (Salaries + Rent + Supplies + Foregone Interest) = 20000 - (5000 + 600 + 1000 + 100)

= 20000 - 6700 = 13300

Answer is

$13,400; $13,300.